OpenStudy (anonymous):

Find an equation of the tangent line at the point specified. y= e^t (1- cos 18t ), t= pi/36

OpenStudy (badhi):

when $$t=\frac\pi{36}$$ $$y=e^{\frac\pi{36}}\left(1-\cos 18\frac\pi{36}\right)=e^{\frac\pi{36}}\left(1-\cos \frac\pi{2}\right)=e^{\frac\pi{36}}$$ so it is clear that the line goes through the point $$(x_0,y_0)=\left(\frac{\pi}{36},e^\frac\pi{36}\right)$$ you can find the slope of the line by differentiating y wrt t, (m) find the equation of the line by, $$\frac{y-y_0}{x-x_0}=m$$

OpenStudy (anonymous):

thanks.

OpenStudy (badhi):

youre welcome :)